Find the ordered pair solution to the following system of equations (Use any method):
3x – y = 5
2x + 5y = -8

Answer by mananth(16946)

3x – y = 5
2x + 5y = -8
..
multiply equation 1 by 5 and add it to equation 2
15x-5y=25
2x+5y=-8
..
17x=17
x=1
..
plug value of x in equation 1
3-y=5
y=-2